The following data is provided by Cotton Tops, Inc a distributor of budget-priced, custom designed tee-shirts that it sells to concessionaires at various theme parks in the US. Use the following information to solve for the Economic Ordering Quantity (EOQ) and how many times a year should Cotton Tops, Inc place an order. The annual sale is 26,000 shirts per year Carrying costs are 25% of inventory value Purchase price per shirt is $4.92 Fixed cost per order is $1,000 (mainly from the labor cost of setting up the equipment)
Answer #1

Ans.EOQ = 2RO/C

Where EOQ stands for Economic Order Quantity

R= Annual Requirement/Demand

O= Order Cost

C= Carrying Cost

R= 26000 units    O = $ 1000 per order          C = 25% of inventory value    

Purchase price = $ 4.92   Therefore, C = 0.25*4.92 = $1.23

Therefore EOQ = 2 * 26000 * 1000/1.23 = 6502 units

Number of Order in a year = R/EOQ = 26000/6502 = 3.99 = 4 Order (Approx.) every year

Number of times an order is placed throughout a year = 12 month/4 = 3 month i.e after every 3 months an order is placed.